Use the graph to find the following.
(a) the domain of f
(b) the range of f  
(c) the x-intercepts
(d) the y-intercept
(e) intervals on which f is increasing
(f) intervals on which f is decreasing
(g) intervals on which f is constant
(h) the number at which f has a relative minimum
(i) the relative minimum of f   
(j) f(0)(k) The values of x for which f(x)=3(l) Is f even, odd or neither?
The 1st coordinate is -6,24. The 2nd coordinate is 0,24.The 3rd coordinate is 5,-1.The 4th coordinate is 10,26.
1
Expert's answer
2020-09-01T17:32:06-0400


The function is


"f(x)= \begin{cases}
 24 & x<0 \\
 ax^2+bx+c &x\geq0 
\end{cases}"

"f(0)=24: 24=c"

"f(5)=-1: -1=a(5)^2+b(5)+24"

"f(10)=26: 26=a(10)^2+b(10)+24"


"\begin{matrix}
 25a+5b=-25 \\
 100a+10b=2 
\end{matrix}"

"\begin{matrix}
 b=-5a-5 \\
 25a=26 
\end{matrix}"


"\begin{matrix}
 a=1.04 \\
 b=-10.2 
\end{matrix}"


"f(x)= \begin{cases}
 24 & x<0 \\
 1.04x^2-10.2x+c &x\geq0 
\end{cases}"

"x_{vertex}=-\dfrac{b}{2a}=-\dfrac{-10.2}{2(1.04)}=\dfrac{255}{52}"

"y_{vertex}=1.04(\dfrac{255}{52})^2-10.2(\dfrac{255}{52})+24=-\dfrac{105}{104}"


"f(0)=24"


"y=0: 1.04x^2-10.2x+24=0"

"1.04x^2-10.2x+24=0"

"26x^2-255x+600=0"

"x=\dfrac{255\pm\sqrt{(-255)^2-4(26)(600)}}{2(26)}=\dfrac{255\pm5\sqrt{105}}{52}"

"x_1=\dfrac{255-5\sqrt{105}}{52}, x_2=\dfrac{255+5\sqrt{105}}{52}"

"f(x)=3=1.04x^2-10.2x+24"


"1.04x^2-10.2x+21=0"

"26x^2-255x+525=0"

"x=\dfrac{255\pm\sqrt{(-255)^2-4(26)(525)}}{2(26)}=\dfrac{255\pm5\sqrt{417}}{52}"

"x_1=\dfrac{255-5\sqrt{417}}{52}, x_2=\dfrac{255+5\sqrt{417}}{52}"

(a) Domain: "(-\infin,\infin)"


(b) Range: "( -\dfrac{105}{104}, \infin)"


(c) the x-intercepts

"(\dfrac{255-5\sqrt{105}}{52}, 0), (\dfrac{255+5\sqrt{105}}{52}, 0)"


(d) the y-intercept

"(0,24)"


(e) intervals on which f is increasing

"(\dfrac{255}{52}, \infin)"


(f) intervals on which f is decreasing

"(0, \dfrac{255}{52})"


(g) intervals on which f is constant

"(-\infin, 0)"


(h) the number at which f has a relative minimum

"\dfrac{255}{52}"


(i) the relative minimum of f

"-\dfrac{105}{104}"


(j) f(0)

"24"


(k) The values of x for which f(x)=3

"\dfrac{255-5\sqrt{417}}{52}, \dfrac{255+5\sqrt{417}}{52}"


(l) Is f even, odd or neither?

The function f is neither even nor odd.